# MITO: Enabling Non-Line-of-Sight Perception using Millimeter-waves through Real-World Datasets and Simulation Tools

### v1

@@ 8.1 Non-Line-of-Sight Perception Tasks
MITO paves the way for future work in multiple directions:
mmWave Object Segmentation & Classification. While we presented initial approaches for mmWave segmentation and classification, future work can improve these models to achieve higher accuracy. This will enable many new applications such as confirming customer orders are properly packaged in e-commerce warehouses, detecting broken objects during shipping, and other applications.
@@
Object Completion. In some cases, the mmWave images do not capture the entire object (See supplementary for more details). This may be due to object specularity, where signals are reflected away from the radar, and therefore parts of the object will not be imaged. It would be interesting future work to create an object completion network that learns the missing parts of an object & creates more photo-realistic NLOS images.
@@
6DoF Estimation. Another interesting problem is 6DoF estimation from mmWave images. This could be used to ensure proper liquid packaging to avoid spills, enable efficient grasp planning in NLOS (e.g., grasping an item under cloth or from within a box filled with packing peanuts), etc.
@@
Robotic Manipulation. Future work can leverage many of the above NLOS perception tasks to enable new tasks in robotic manipulation. For example, accurate mmWave object segmentation and classification can signficiantly increase the efficiency of mechanical search and retrieval (i.e., the problem of searching for a specific item in clutter). Further, mmWave images can be used to analyze pile stability to improve push and grasp planners. NLOS perception can also help autonomous robots determine which packages contain fragile materials to be handled with special care. More generally, NLOS perception can help autonomous robots to operate in more realistic, cluttered environments.
8.2 Dataset Extensions.
We will continue to extend MITO’s dataset and simulation in various ways, including: (1) scenarios with multiple items, (2) other surfaces (e.g., tables) on which the items are placed, (3) additional synthetic data, e.g., 3D meshes from existing datasets in addition to real-world images here, (4) capturing objects from more angles. In addition, it would be helpful to advance the efficiency of data capture in the real-world experiments, and to speed-up our simulator. [^9] Our simulator currently takes a median of 2360 seconds to produce all synthetic signals for an object (Eqs. 4/5) and 21 seconds to produce an image (Eq. 1) when run on a computer with a a Intel Core i9-10900X CPU and one GeForce RTX 2080 Ti GPU.
@@ 9 Conclusion
We presented MITO, the first multi-spectral mmWave dataset for everyday objects in both line-of-sight and fully-occluded settings. We provide over 580 real world images from 76 (out of 80) objects in the YCB dataset, as well as an open-source simulation tool that can be used to generate synthetic images for any 3D mesh. We showed two applications of MITO: object segmentation and classification. As we expand MITO with real and synthetic images, we hope that it enables the broader community to create the same advancements in mmWave non-line-of-sight perception that other datasets enabled for RGB perception.

### v3

@@ 6.3 Additional NLOS Applications & Extensions
In addition to the applications described above, MITO also paves the way for future work in multiple directions:
@@
  • NLOS Object Completion: In some cases, the mmWave images do not capture the entire object due to specularity (See supplementary for more details). It would be interesting future work to create an object completion network that learns the missing parts of an object & creates more photo-realistic NLOS images.
@@
  • NLOS 6DoF Estimation: Another interesting problem is 6DoF estimation from mmWave images. This could be used to ensure proper liquid packaging to avoid spills, or enable efficient grasp planning in NLOS (e.g., grasping an item under cloth or within a box with packing peanuts).
@@
  • NLOS Robotic Manipulation: Future work can leverage many of the above NLOS perception tasks to enable new tasks in robotic manipulation. For example, accurate mmWave object segmentation and classification can significantly increase the efficiency of mechanical search and retrieval (i.e., the problem of searching for a specific item in clutter). Further, mmWave images can be used to analyze pile stability to improve push and grasp planners. NLOS perception can also help autonomous robots determine which packages contain fragile materials to be handled with special care. More generally, NLOS perception can help autonomous robots to operate in more realistic, cluttered environments.
@@ 7 Conclusion
We present MITO, the first millimeter-wave (mmWave) dataset of diverse, everyday objects in both line-of-sight and fully-occluded settings. We collect over 24 million frames of mmWave images, and used them to synthesize 550 high-resolution real world 3D (SAR) mmWave images for 76 objects. We also develop an open-source simulator that can be used to generate synthetic images for any 3D mesh. We demonstrate the utility of our contributions in two applications – NLOS object segmentation and NLOS shape classification – establishing baselines for future research and paving the way for important applications in robotic perception, logistics, scene understanding.
As the research evolves, it would be interesting to expand the dataset – incorporating additional items, surfaces, occlusion types, etc. – which would continue improving the accuracy of the demonstrated NLOS tasks and unlock additional ones. More generally, we hope this work serves as a foundation and motivation for future research in mmWave-based NLOS perception, similar to how early RGB datasets accelerated the research in optical-based Computer Vision, and continued driving the field forward as they expanded.